# NEDS-SLAM: A Novel Neural Explicit Dense Semantic SLAM Framework using 3D Gaussian Splatting

### v2

@@ Abstract
We propose NEDS-SLAM, an Explicit Dense semantic SLAM system based on 3D Gaussian representation, that enables robust 3D semantic mapping, accurate camera tracking, and high-quality rendering in real-time. In the system, we propose a Spatially Consistent Feature Fusion model to reduce the effect of erroneous estimates from pre-trained segmentation head on semantic reconstruction, achieving robust 3D semantic Gaussian mapping. Additionally, we employ a lightweight encoder-decoder to compress the high-dimensional semantic features into a compact 3D Gaussian representation, mitigating the burden of excessive memory consumption. Furthermore, we leverage the advantage of 3D Gaussian splatting, which enables efficient and differentiable novel view rendering, and propose a Virtual Camera View Pruning method to eliminate outlier GS points, thereby effectively enhancing the quality of scene representations. Our NEDS-SLAM method demonstrates competitive performance over existing dense semantic SLAM methods in terms of mapping and tracking accuracy on Replica and ScanNet datasets, while also showing excellent capabilities in 3D dense semantic mapping.
@@ I Introduction
Inspired by the success of NERF and 3D Gaussian Splatting (3DGS) in high-fidelity view synthesis, researchers have explored building end-to-end visual SLAM systems based on neural radiance fields. These novel SLAM architectures offer superior solutions compared to traditional algorithms in terms of surface continuity, memory requirements, and scene completion. Specifically, iMAP [2] and NICE-SLAM [3] leverage neural implicit fields for consistent geometry representation, while MonoSLAM and SplatAM employ 3D Gaussian Splatting (3DGS) to achieve photo-realistic mapping.
Existing radiance field-based SLAM methods primarily focus on RGB reconstruction, while only a few approaches, such as SGS-SLAM [4] and DNS-SLAM [5], address semantic reconstruction. However, these methods rely on accurate and consistent semantic pre-segmentation, and their reconstruction performance suffers significantly from any substantial deviations in semantic feature estimation. In practical SLAM scenarios, the inconsistency of semantic feature estimation is a non-negligible issue.
Overall, Neural Explicit Dense Semantic SLAM can be summarized as facing two key challenges: 1) Providing robust semantic reconstruction results under inconsistent semantic features. 2) Incrementally building a map that can accurately distinguish well-optimized and low-quality regions, while effectively filtering out outliers to improve reconstruction quality.
@@
  • We propose a fusion module that combines semantic features with appearance features, addressing the spatially inconsistency of semantic features and providing a more robust semantic SLAM solution.
  • We build a semantic SLAM framework based on 3DGS, embedding semantic features with a lightweight encoder-decoder to achieve accurate semantic reconstruction and photo-realistic reconstruction.
  • We introduce a virtual camera view pruning method to remove noisy Gaussians, enabling more accurate construction of the 3D Gaussian radiance field.
@@ II-A Traditional approaches to dense semantic SLAM
Semantic information is paramount for SLAM systems, as it enables scene understanding beyond geometric reconstruction alone. The ability to perceive and model semantics represents a crucial requirement for enabling applications of SLAM in robotics, virtual or augmented reality, and other domains that require understanding of the environment [6],[7]. Real-time dense semantic SLAM systems face the challenge of effectively fusing semantic information into underlying 3D geometric representations of the environment. Traditional approaches using voxels, point clouds, and signed distance fields to encode object labels [8],[9]. However, voxel- and point cloud-based approaches struggle with reconstruction speed and high-fidelity model acquisition. Meanwhile, signed distance field representations incur high memory usage that does not scale well to large-scale environments. There remains a need for more efficient and expressive 3D semantic modeling techniques suitable for real-time dense SLAM.
II-B Gaussian Splatting based SLAM
3D Gaussian representations have emerged as a promising approach for 3D scene modelling using a set of 3D Gaussians, each characterized by parameters such as position, anisotropic covariance, opacity, and color [10]. While existing GS-based SLAM methods have primarily focused on RGB reconstruction, exploring end-to-end system architectures, optimization of GS parameters, and accurate camera pose tracking through differentiable rendering, less attention has been paid to semantic reconstruction [11],[12],[13],[14] The few semantic GS SLAM approaches proposed to date have simply encoded ground truth semantic color labels directly as a second color channel of the GS parameters [4], without explicit modeling of semantic information or inference. There is clear potential for more sophisticated integration of semantics within the GS framework. The present work conducts a more in-depth exploration of semantic GS SLAM, aiming to simultaneously improve the robustness and reconstruction fidelity of GS-based SLAM systems through more sophisticated modeling and inference of semantic information within the GS representation.
@@ III-A Scene Representation and Semantic embedding
Each 3DGS utilized for representing three-dimensional scenes encompasses mean, covariance, and color information. In this paper, a simplified 3DGS representation of the scene is employed [11], omitting the spherical harmonics functions used for color representation, while assuming GS to be isotropic.
$f^{gs}\left(\mathbf{x}\right)=o\exp\left({-\frac{{{{\left\|{\mathbf{x}-\mu}\right\|}^{2}}}}{{2{r^{2}}}}}\right)$ (1)
Where $\mu\in\mathbb{R}^{3}$ represents the center position of the GS, $r$ is the radius, and $o\in[0,1]$ represents the opacity. The rapid and differentiable rendering based on GS splatting serves as the core of mapping and tracking within GS-based SLAM systems. This ability for fast rendering enables the system to directly compute the gradients of the underlying GS parameters based on the discrepancy between the rendered results and the actual data. Consequently, the GS parameters can be updated to achieve an accurate representation of the scene. The differentiable rendering process based on GS splatting comprises three steps: Frustum Culling, Splatting, and Rendering by Pixels [15].
$C\left(p\right)=\sum\limits_{i\in N}{\mathbf{c}_{i}}f_{i}^{gs}\left(p\right)\prod\limits_{j=1}^{i-1}{\left({1-f_{j}^{gs}\left(p\right)}\right)}$ (2)
After arranging a collection of 3D Gaussians and camera pose, it is imperative to sort the Gaussians in a front-to-back manner. By employing alpha-compositing, the splatted 2D projection of each Gaussian can be efficiently rendered in pixel space, ensuring the generation of RGB images in the desired order, as Eq. (2). $\mathbf{c}_{i}$ represents the color parameters of the GS, and $f_{i}^{gs}\left(p\right)$ is computed as in Eq. (1) but with the 2D splatted ${\mu}$ and ${r}$. The rendering process is completed by multiplying the opacity of each GS with the color and accumulating the results. The depth map is rendered in a similar manner, as shown in Eq. (3).
$D\left(p\right)=\sum\limits_{i\in N}{\mathbf{d}_{i}}f_{i}^{gs}\left(p\right)\prod\limits_{j=1}^{i-1}{\left({1-f_{j}^{gs}\left(p\right)}\right)}$ (3)
@@
A simple way to combine 3D Gaussian Splatting with semantic features is to add trainable feature vectors to each Gaussian distribution. These parameters can be learned during the differentiable rendering process, which allows end-to-end training. However, for explicit Dense Semantic SLAM, adding a high dimensional semantic feature vector to each 3DGS is memory-inefficient. It would also weaken the real-time performance of optimizing model parameters. Inspired by LangSplat [17], we propose using a simple MLP as an encoder to compact high-dimensional semantic features into a low-dimensional vector. The compressed semantic features are then added to the 3D gaussian splats and can be rendered as in Eq. (4).
$S\left(p\right)=\sum\limits_{i\in N}{\mathbf{f}_{i}}f_{i}^{gs}\left(p\right)\prod\limits_{j=1}^{i-1}{\left({1-f_{j}^{gs}\left(p\right)}\right)}$ (4)
@@ III-B Adaptive 3D Gaussian Expansion Mapping
III-B1 Spatially Consistent Feature Fusion
@@
Consistent and continuous semantic labels are crucial for 3D semantic mapping. Semantic SLAM uses pretrained semantic segmentation models to compute pixel-level semantic labels from each RGB frame, but these class labels lack environmental specificity. Pretrained models may produce inconsistent semantic estimates, where the same object is predicted with different semantic labels in images from different angles. This would significantly reduce the quality of constructing 3D semantic Gaussians.
Figure 2: Using Mask2Former [18] for semantic segmentation on the Replica dataset, there is a noticeable inconsistency in semantic labels between frame 0 and frame 30 (highlighted in white boxes in the image). The floor and chairs exhibit significant differences, and directly utilizing the semantic distribution of the scene for GS construction would result in incompact semantics.
@@
To address this issue, SNI-SLAM [19] computes a fused feature by combining geometry, appearance, and semantic features, replacing the reliance on a single semantic label. CoSSegGaussians [20] incorporates DINO [16] features with superior multi-view semantic scale consistency into the GS parameters. Subsequently, the semantic encoding of each GS is fused with spatial coordinates to render semantic features, thereby enhancing robustness. Inspired by SNI-SLAM and CoSSegGaussians, this paper proposes a simplified fusion mechanism. It combines the appearance features estimated by DepthAnything [21] with the semantic features extracted from pretrained model. The resulting mixed feature, obtained through MLP computations, is then embedded as the final semantic encoding in the 3DGS representations.
Due to limitations imposed by sensor performance and real-world noise, the depth map is not densely populated at the pixel level. The utilization of DepthAnything as spatial feature extractor to estimate the relative depth map of the current frame aims to address the issue of holes and artifacts in the depth observation. Within the SLAM pipeline, an effective mask is computed for each depth frame, rendering pixels without depth information as masked out. These masked pixels are unable to contribute to the establishment and optimization of subsequent GS, even if they possess accurate RGB and semantic values. By employing DepthAnything, it ensures that the geometric information of the entire scene within the field of view is incorporated into the extraction of semantic features and the modeling process of semantic GS.
@@
The relative depth between pixels can reflect the geometric structure of observed object surfaces. The feature fusion model dynamically adjusts the weights of semantic features according to the spatially consistent relative relationships between objects. It thereby reduces the impact of segmentation errors on the spatial consistency of semantic features.
@@ III-B2 Updating 3D Gaussians
During the mapping process, we assume that the camera pose for the current frame is known. We need to use the current keyframe’s RGBD data to update the GS model of the scene. The updating has two meanings: one is to optimize the established scene parameters, and the other is to generate the new explored GS distribution of the scene.
Following the processes used in Splatam [11] and GS-SLAM [12], silhouette images are rendered to determine the contribution of each surface element (GS) to the map. At the same time, the difference between the projected depth value and the ground truth value of pixels corresponding to newly added GS is checked when they are projected back onto the image plane.
@@
By setting ${\mathbf{d}_{i}}$ in Eq . 3 to a unit value, the silhouette of pixel $p$ is computed. The densification mask $M\left(p\right)$ is calculated according to Eq . 5, where $D\left(p\right)$ represents the depth value of pixel $p$. By calculating the difference between the rendered output from a specific camera position and the ground truth value using Gaussian splatting, the parameters of the 3D Gaussian distribution can be optimized. This problem can also be described as fitting an explicit radiance field to images where the camera pose is known.
After the modeling process discussed in Section III-A, the modeled scene contains three feature channels: spatial position, surface color, and potential semantics. The spatial position and surface color are directly obtained from the RGBD data stream. Meanwhile, the fusion of semantic encoding is supervised by the mask output from a pretrained segmentation model.
@@
The color loss $L_{c}$ is represented as a weighted combination of SSIM [10] and $L1$ loss as in Eq. 6.
@@
The depth loss $L_{d}$ is calculated as in Eq. 7. During the mapping stage, the multi-channel loss is as shown in Eq. 8, where $S_{render}$ represents the rendered semantic feature and $S_{head}$ represents the segmentation mask of the current frame computed by the pretrained model.
${L_{mapping}}={\lambda_{c}}{L_{c}}+{\lambda_{d}}{L_{d}}+{\lambda_{s}}{L_{1}}\left({{S_{render}},{S_{head}}}\right)$ (8)
@@
In Eq. 8, $\lambda_{d}$, $\lambda_{s}$, and $\lambda_{c}$ are predefined hyperparameters used to assign weighted values to the depth, semantic, and color channels respectively.
III-B3 Vitrual Camera Pruning 3D Gaussians

### v3

# NEDS-SLAM: A Neural Explicit Dense Semantic SLAM Framework using 3D Gaussian Splatting

@@ Abstract
We propose NEDS-SLAM, a dense semantic SLAM system based on 3D Gaussian representation, that enables robust 3D semantic mapping, accurate camera tracking, and high-quality rendering in real-time. In the system, we propose a Spatially Consistent Feature Fusion model to reduce the effect of erroneous estimates from pre-trained segmentation head on semantic reconstruction, achieving robust 3D semantic Gaussian mapping. Additionally, we employ a lightweight encoder-decoder to compress the high-dimensional semantic features into a compact 3D Gaussian representation, mitigating the burden of excessive memory consumption. Furthermore, we leverage the advantage of 3D Gaussian splatting, which enables efficient and differentiable novel view rendering, and propose a Virtual Camera View Pruning method to eliminate outlier gaussians, thereby effectively enhancing the quality of scene representations. Our NEDS-SLAM method demonstrates competitive performance over existing dense semantic SLAM methods in terms of mapping and tracking accuracy on Replica and ScanNet datasets, while also showing excellent capabilities in 3D dense semantic mapping.
@@ I Introduction
Inspired by the success of NERF and 3D Gaussian Splatting (3DGS) in high-fidelity view synthesis, researchers have explored building end-to-end visual SLAM systems based on neural radiance fields. These novel SLAM architectures offer superior solutions compared to traditional algorithms in terms of surface continuity, memory requirements, and scene completion. Specifically, iMAP [2] and NICE-SLAM [3] leverage neural implicit fields for consistent geometry representation, while MonoGS [4] and SplaTAM [5] employ 3DGS to achieve photo-realistic mapping.
Given continuous input of RGB-D frames, dense semantic SLAM aims to create a compact and dense 3D representation of the scene that includes accurate RGB information as well as dense semantic data. However, current state-of-the-art semantic segmentation models are trained on large amounts of internet images, which are loosely related and time-independent. This leads to estimation errors such as semantic spatial inconsistency, which significantly impairs the density and completeness of semantic reconstruction. The previous 3DGS-based semantic SLAM method [6] overlooked the issue of semantic feature inconsistency, which limits its potential for practical applications.
Furthermore, our research has found that directly embedding semantic category labels into gaussians parameters may not be appropriate. During splatting, overlapping gaussians combine through alpha-blending to form pixel values on the imaging plane. Using RGB color channels as an example, ideally, when 3D gaussians are splatted onto different imaging planes, they create different color blending effects. However, assigning fixed class labels to the gaussians leads to meaningless values in the semantic channels during splatting. Therefore, attempting to embed semantic features instead of semantic category labels into the 3D gaussians parameters would be more promising. However, this approach can cause prohibitive memory requirements and significantly lower the efficiency of both optimization and rendering, as semantic features typically have higher dimensions, whereas category labels are just integer values.
In a 3DGS-based SLAM system, the process of incrementally building a map is often influenced by camera pose estimation errors, object occlusions, and errors in the optimization process. These factors can introduce 3D gaussians that do not align with actual surfaces. When these outlier gaussians are included in the rendering view, they can create visual artifacts, which in turn affect camera pose estimation, creating a vicious cycle. This issue is not addressed in the original 3DGS paper, where the camera poses for each frame are precomputed using an offline SFM method. Therefore, handling outlier gaussians is crucial for 3DGS-based SLAM methods.
Overall, 3DGS based Dense Semantic SLAM can be summarized as facing two key challenges: 1) Providing robust semantic reconstruction results under inconsistent semantic features. 2) Incrementally building a map that can accurately distinguish well-optimized and low-quality regions, while effectively filtering out outliers to improve reconstruction quality.
@@
  • We propose the Spatially Consistent Feature Fusion module (SCFF), which combines semantic features with appearance features. This module addresses the spatial inconsistency of semantic features and provides a more robust semantic SLAM solution.
  • We embed semantic features into Gaussian parameters instead of using category labels. We also introduce a lightweight encoder-decoder to prevent memory issues from high-dimensional semantic feature embedding.
  • We present the Virtual Camera View Pruning (VCVP) method. VCVP generates multiple virtual camera views to ensure consistency, identifying and removing unstable gaussians caused by occlusions, camera pose errors, and parameter optimization issues, leading to a more accurate 3D Gaussian field.
@@ II-A Traditional approaches to dense semantic SLAM
Real-time dense semantic SLAM systems face the challenge of effectively fusing semantic information into underlying 3D geometric representations of the environment. Traditional approaches use voxels, point clouds, and signed distance fields to encode object labels [7],[8]. However, voxel- and point cloud-based approaches struggle with reconstruction speed and high-fidelity model acquisition. Meanwhile, signed distance field representations incur high memory usage that does not scale well to large-scale environments. There remains a need for more efficient and expressive 3D semantic modeling techniques suitable for real-time dense SLAM.
II-B NeRF based SLAM
In recent years, Neural Radiance Fields (NeRF) have sparked significant interest in computer graphics, attracting attention for their high-fidelity novel view synthesis and lightweight scene representation [9]. This enthusiasm has quickly spread to the SLAM field, leading to the development of many innovative SLAM architectures [2][3]. Zhu et al. introduced SNI-SLAM [10], which employs neural implicit representation and hierarchical semantic encoding for multi-level scene understanding, contributing a cross-attention mechanism for the collaborative integration of appearance, geometry, and semantic features. Due to the limitations of NeRF’s volume rendering, NeRF-based dense semantic SLAM struggles to simultaneously model and optimize the semantic and RGB-geometry information of the environment [11][12]. Additionally, the efficiency of SLAM is constrained by the implicit representation of the map[13].
II-C Gaussian Splatting based SLAM
3DGS representations have emerged as a promising approach for 3D scene modelling using a set of 3D gaussians, each characterized by parameters such as position, anisotropic covariance, opacity, and color [14]. While existing 3DGS-based SLAM methods have primarily focused on RGB reconstruction, exploring end-to-end system architectures, optimization of gaussians parameters, and accurate camera pose tracking through differentiable rendering, less attention has been paid to semantic reconstruction [5],[15],[4],[16]. The few semantic 3DGS-SLAM approaches proposed to date have simply encoded ground truth semantic color labels directly as a second color channel of the gaussians parameters [6], without explicit modeling of semantic information or inference. There is clear potential for more sophisticated integration of semantics within the 3DGS-SLAM framework. The present work conducts a more in-depth exploration of dense semantic SLAM, aiming to simultaneously improve the robustness and reconstruction fidelity of 3DGS-based SLAM systems through more sophisticated modeling and inference of semantic information within the 3DGS representation.
@@ III-A Scene Representation and Semantic embedding
Each 3DGS utilized for representing three-dimensional scenes encompasses mean, covariance, and color information. In this paper, a simplified 3DGS representation of the scene is employed [5], omitting the spherical harmonics functions used for color representation, while assuming gaussians to be isotropic as in Eq 1.
$f^{gs}\left(\bm{x}\right)=o\exp\left({-\frac{{{{\left\|{\bm{x}-\mu}\right\|}^{2}}}}{{2{r^{2}}}}}\right)$ (1)
Where $\mu\in\mathbb{R}^{3}$ represents the center position, $r$ is the radius, and $o\in[0,1]$ represents the opacity. The rapid and differentiable rendering based on 3DGS serves as the core of mapping and tracking within 3DGS-based SLAM systems. This ability for fast rendering enables the system to directly compute the gradients of the underlying parameters based on the discrepancy between the rendered results and the actual data. Consequently, the gaussians parameters can be updated to achieve an accurate representation of the scene. The differentiable rendering process based on gaussians splatting comprises three steps: Frustum Culling, Splatting, and Rendering by Pixels [17].
$C\left(p\right)=\sum\limits_{i\in N}{\bm{c}_{i}}f_{i}^{gs}\left(p\right)\prod\limits_{j=1}^{i-1}{\left({1-f_{j}^{gs}\left(p\right)}\right)}$ (2)
After arranging a collection of 3D gaussians and camera pose, it is imperative to sort the gaussians in a front-to-back manner. By employing alpha-compositing, the splatted 2D projection of each gaussian can be efficiently rendered in pixel space, ensuring the generation of RGB images in the desired order, as Eq 2. $\bm{c}_{i}$ represents the color parameters of the gaussians, and $f_{i}^{gs}\left(p\right)$ is computed as in Eq 1 but with the 2D splatted ${\mu}$ and ${r}$. The rendering process is completed by multiplying the opacity of each gaussian with the color and accumulating the results. The depth map is rendered in a similar manner, as shown in Eq 3.
$D\left(p\right)=\sum\limits_{i\in N}{\bm{d}_{i}}f_{i}^{gs}\left(p\right)\prod\limits_{j=1}^{i-1}{\left({1-f_{j}^{gs}\left(p\right)}\right)}$ (3)
@@
A simple way to combine 3DGS with semantic features is to add trainable feature vectors to each gaussian. These parameters can be learned during the differentiable rendering process, which allows end-to-end training. However, for dense semantic SLAM, adding a high dimensional semantic feature vector to each 3DGS is memory-inefficient. Inspired by LangSplat [19], we propose using a simple MLP as an encoder to compact semantic features into a low-dimensional vector. The compressed semantic features are then added to the 3D gaussians and can be rendered as in Eq 4.
$S\left(p\right)=\sum\limits_{i\in N}{\bm{f}_{i}}f_{i}^{gs}\left(p\right)\prod\limits_{j=1}^{i-1}{\left({1-f_{j}^{gs}\left(p\right)}\right)}$ (4)
@@ III-B Adaptive 3D Gaussian Expansion Mapping
III-B1 Spatially Consistent Feature Fusion (SCFF)
@@
Pervious semantic SLAM approaches typically use pretrained segmentation models to compute pixel-level labels from each RGB frame,but these class labels lack environmental specificity. Pretrained models may produce inconsistent semantic estimates, where the same object is predicted with different semantic labels in images from different camera views.
@@
To address this issue, SNI-SLAM [10] computes a fused feature by combining geometry, appearance, and semantic features. CoSSegGaussians [20] incorporates DINO [18] features with superior multi-view semantic scale consistency into the gaussians parameters. Subsequently, the semantic encoding of each gaussians is fused with spatial coordinates to render semantic features, thereby enhancing robustness.
In this paper, we propose a simplified fusion mechanism. It combines the appearance features with the semantic features extracted from pretrained model. The resulting mixed feature, obtained through MLP encoder, is then embedded as the final semantic encoding in the 3DGS representations.
As shown in Fig 1, the pretrained semantic feature extractor extracts an $H\times W\times D_{f}$ feature map $F_{df}$ from an $H\times W\times 3$ RGB frame. At the same time, the spatial feature extractor extracts $H\times W\times D_{s}$ features $F_{ds}$ from RGB data. After three layers of convolution, the feature channels of $F_{df}$ are reduced to 256, 128, and 16, respectively. Similarly, the feature channels of $F_{ds}$ are increased to 16 through one layer of CNN. After concatenation and the final convolution, we obtain the spatially consistent feature $F^{32}_{scff}$ with 32 channels. Using the external parameters of the camera, we can convert an input frame of RGBD into a series of points in 3D space. Each point includes $xyz$ coordinates, RGB information, and 32-channel SCFF features.
To reduce the number of 3D gaussian parameters, we need to use an information encoding method to compress $F^{32}_{scff}$ to a lower dimension, such as using hash encoding [21], GPR [22], etc. In this paper, we use a simple MLP to compress $F^{32}_{scff}$ to three dimensions, resulting in $F^{3}_{scff}$.
It is important to clarify that the semantic category labels are numerical IDs from a predefined category library (e.g., 0 represents a person), while the $F^{3}_{scff}$ values range between 0 and 1. These semantic features can be decoded back into semantic category labels by a subsequent decoder.
We use the pre-trained DINO [18] model as a semantic feature extractor, obtaining features $F_{df}$ with 384 channels ($D_{f}=384$). We use DepthAnything [23] as the spatial feature extractor, resulting in $D_{s}=1$. Relative depth output from DepthAnything is used as the appearance feature because changes in the camera viewpoint do not affect the relative position of surfaces on the object. The spatial consistency of appearance features helps SCFF achieve stable semantic feature estimation.
@@
The relative depth between pixels can reflect the geometric structure of observed surfaces. The SCFF module dynamically adjusts the weights of semantic features according to the spatially consistent relationships. It thereby reduces the impact of segmentation errors on the spatial consistency of semantic features.
@@ III-B2 Updating 3D Gaussians
During the mapping process, we assume that the camera pose for the current frame is known. We need to use the current keyframe’s RGBD data to update the gaussians representation of the scene. Updating has two meanings: optimizing existing scene parameters and generating a new 3DGS distribution for the scene.
Following the processes used in Splatam [5] and GS-SLAM [15], we use Eq.5 to calculate the silhouette value per pixel. The silhouette images are rendered to determine the contribution of each gaussian to the map.
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}Sil\left(p\right)=\sum\limits_{i\in N}f_{i}^{gs}\left(p\right)\prod\limits_{j=1}^{i-1}{\left({1-f_{j}^{gs}\left(p\right)}\right)}}$ (5)
At the same time, the difference between the projected depth value and the ground truth value of pixels corresponding to newly added gaussian is checked when they are projected back onto the image plane.
@@
The densification mask $M\left(p\right)$ is calculated according to Eq 6, where $D\left(p\right)$ represents the depth value of pixel $p$. $M\left(p\right)$ represents a Boolean mask for pixel $p$. The optimization of 3DGS and the addition of new gaussians will be confined to areas where the mask value is True, thereby avoiding the densification of gaussians in well-reconstructed areas. This differs from the approach in [14], which splits gaussians in over-reconstructed regions. Due to the high real-time requirements of SLAM systems, setting threshold parameters in $M\left(p\right)$ allows the system to avoid the heavy computation associated with the gaussians densification method in [14].
After the process discussed in Section III-A, the scene representations contains three feature channels: spatial position, surface color, and potential semantics. The spatial position and surface color are directly obtained from the RGBD data stream. Meanwhile, the fusion of semantic encoding is supervised by the mask output from a pretrained segmentation model.
@@
The color loss $L_{c}$ is represented as a weighted combination of SSIM [14] and $L1$ loss as in Eq 7.
@@
The depth loss $L_{d}$ is calculated as in Eq 8. During the mapping stage, the multi-channel loss is as shown in Eq 9, where $S_{render}$ represents the semantic labels after decoding the semantic features and $S_{head}$ represents the class labels computed by the pretrained model. We use the cross-entropy loss ${L_{CE}}$ to supervise the semantic channel.
${L_{mapping}}={\lambda_{c}}{L_{c}}+{\lambda_{d}}{L_{d}}+{\lambda_{s}}{L_{CE}}\left({{S_{render}},{S_{head}}}\right)$ (9)
@@
In Eq 9, $\lambda_{d}$, $\lambda_{s}$, and $\lambda_{c}$ are predefined hyperparameters used to assign weighted values to the depth, semantic, and color channels respectively.
III-B3 Vitrual Camera View Pruning 3D Gaussians (VCVP)